OpenStudy (anonymous):

Two angles are supplementary. One angle measures 26 degrees more than the other. Find the measure of the two angles.

OpenStudy (anonymous):

suuplementary means the two angles add up to 180 so x+y = 180 one angle is 26 more than the other so y = x+26 solve the two equations

OpenStudy (anonymous):

supplementary means their sum is 180 let smaller angle be x deg then bigger angle is (x + 26) deg x + x + 26 = 180 2x = 180-26 x = 154/2 x = 77 so the smaller angle is 77 deg bigger angle is 103 deg

OpenStudy (anonymous):

Supplementaries add up to 180. So take two 90 degree angles, divide your difference in two (13) now subtract that from one angle, add to the other. 90 - 13 = 77, 90 +13 = 103. 103 - 77 = 26 103 + 77 = 180 (supplementary)

OpenStudy (anonymous):

Basically, you are moving your vertical 90-90 line the amount of degrees difference one way or the other. This way, you can do it without algebra.